OpenStudy (anonymous):

Solve for x: −3|2x + 6| = −12 A) x = 1 and x = 5 B) x = −1 and x = −5 C) x = −9 and x = 3 D) no solution I got x = 5, x = -5.

OpenStudy (solomonzelman):

div both sides by -3

OpenStudy (anonymous):

divide through by -3

OpenStudy (anonymous):

equate left side to plus/minus right side

OpenStudy (anonymous):

and then I have (2x + 6) = -4 im using parenthesis as absolute value signs

OpenStudy (solomonzelman):

hold on! -12 divided by -3 equals what ?

OpenStudy (anonymous):

It would be 2x + 6 = -4 or 2x + 6 = 4

OpenStudy (anonymous):

ohhhh, I forgot 3 was negative! hold on

OpenStudy (anonymous):

-12 divided by -3 is 4

OpenStudy (solomonzelman):

yes

OpenStudy (solomonzelman):

\(\Huge\color{blue}{ \sf |2x + 6| = 4 }\) \(\Huge\color{blue}{ \sf 2x + 6 = 4 }~~~~or~~~~\)\(\Huge\color{blue}{ \sf 2x + 6 = -4 }\) solve each equation

OpenStudy (anonymous):

then I get 5 and negative 5.....like what I had in the first place

OpenStudy (anonymous):

2x = 4-6 = -2

OpenStudy (solomonzelman):

hold on, what do you get for the first equation where the 4 is positive?

OpenStudy (texaschic101):

2x + 6 = 4 2x = 4 - 6 2x = -2 x = -1 2x + 6 = -4 2x = -4 - 6 2x = -10 x = -5

OpenStudy (anonymous):

ohhh, I did that the wrong way around. thank you I get it now